# GeoGraph: Geometric and Graph-based Ensemble Descriptors for Intrinsically Disordered Proteins

### v1

@@ Appendix C Geometric feature calculation
As in [10], we calculate the radius of gyration and asphericity features using the mass-weighted Gyration tensor, $\mathbf{T}\in\mathbb{R}^{3\times 3}$, computed as
@@
Parametrise the power-scaling-law relationship describing how the Euclidean distance between residues scales as a function of their spacing in sequence. Following the implementation used by [32], we fit this relationship to residues spaced at least 5 amino acids apart:
@@
Unlike the other geometric features which are calculated for each conformation separately and then averaged, the Flory scaling parameters are calculated by first averaging the inter-residue distances observed for each spacing across the whole ensemble, then using the optimize.curve_fit function provided by SciPy [37] to fit the ($\nu$, $A_{0}$) parameters.
@@ Appendix D Comparisons with existing IDP models
We evaluate two prominent methods for IDP property prediction: ALBATROSS [24] and STARLING [26]. ALBATROSS is a family of 5 recurrent neural network models, each trained to independently predict one of the ensemble-averaged geometric features $\{R_{e},R_{g},\Delta,\nu,A_{0}\}$ directly from sequence. STARLING is a generative diffusion model which generates a conformational ensemble of IDPs by denoising a latent representation of residue-residue distance maps for each conformation. We follow the method used by [26] for property prediction with STARLING: we sample 1000 conformations using 25 DDIM steps, then using the generated ensemble to calculate the ensemble-averaged geometric feature values for each sequence.
We evaluate the publicly released models for both methods on our test set, however we also note that the IDP datasets used to train ALBATROSS and STARLING notably differ from our training dataset Human–IDRome. In particular, their datasets contain synthetic as well as biological IDP sequences, and the conformational ensembles were generated via coarse-grained MD using an adapted version of the Mpipi force field [19] rather than CALVADOS-2. We therefore additionally retrained these models from scratch on the Human–IDRome dataset, and report results using both the pretrained and retrained versions of these models in Table 1.
@@ D.1 STARLING
Following the preprocessing in STARLING [26], we first downsampled the frames for each sequence in our Human–IDRome dataset to reduce the correlation between conformers. We found that keeping every 20th frame was sufficient to stabilise model training, resulting in 50 conformers for each sequence. After downsampling, we used the same hyperparameters and methodology as in [26] to sequentially retrain the STARLING VAE and DDPM models from scratch using our train and validation splits.
@@ D.2.2 Retraining
We use the same model architecture hyperparameters (number of hidden layers, hidden size) for each feature as used in the published V2 models. We found that we could improve training stability and performance by replacing the loss function used by [24] with the mean of the L1 loss over a batch rather than the sum, and performing a grid search over batch sizes $\{$64, 128, 256}\} and learning rates $\{$1e-3, 5e-3, 1e-2}\} for each model. We report the best test $R^{2}$ score achieved over the grid search for each feature in Table 1, and the hyperparameters used for each model in Table 3.
@@ D.2.3 $R^{2}$ score calculation
The $R^{2}$ scores attained in our evaluation of ALBATROSS are notably lower than those reported in the original ALBATROSS work [24]. This discrepancy can be partly explained by a difference in the definition of $R^{2}$ used between our work and theirs. In [24], the authors define the $R^{2}$ score as the square of the Pearson correlation coefficient between the true and predicted values, whereas here we define $R^{2}$ as the coefficient of determination.
@@ D.2.4 Flory scaling parameters
We compute the Flory scaling parameters by fitting a power-law relationship between the Euclidean distance of residue pairs and their sequence separation. Following the methodology of the Human-IDRome paper [32], we exclude residue pairs with a sequence separation of less than five, as these short-range interactions are governed by local chain stiffness and deviate from the global scaling law, which may differ from that used by ALBATROSS [24].

### v2

@@ Appendix C Geometric feature calculation
As in [dima2004], we calculate the radius of gyration and asphericity features using the mass-weighted Gyration tensor, $\mathbf{T}\in\mathbb{R}^{3\times 3}$, computed as
@@
Parametrise the power-scaling-law relationship describing how the Euclidean distance between residues scales as a function of their spacing in sequence. Following the implementation used by [tesei2024idrome], we fit this relationship to residues spaced at least 5 amino acids apart:
@@
Unlike the other geometric features which are calculated for each conformation separately and then averaged, the Flory scaling parameters are calculated by first averaging the inter-residue distances observed for each spacing across the whole ensemble, then using the optimize.curve_fit function provided by SciPy [2020SciPy-NMeth] to fit the ($\nu$, $A_{0}$) parameters.
@@ Appendix D Comparisons with existing IDP models
We evaluate two prominent methods for IDP property prediction: ALBATROSS [lotthammer2024] and STARLING [novak2025]. ALBATROSS is a family of 5 recurrent neural network models, each trained to independently predict one of the ensemble-averaged geometric features $\{R_{e},R_{g},\Delta,\nu,A_{0}\}$ directly from sequence. STARLING is a generative diffusion model which generates a conformational ensemble of IDPs by denoising a latent representation of residue-residue distance maps for each conformation. We follow the method used by [novak2025] for property prediction with STARLING: we sample 1000 conformations using 25 DDIM steps, then using the generated ensemble to calculate the ensemble-averaged geometric feature values for each sequence.
We evaluate the publicly released models for both methods on our test set, however we also note that the IDP datasets used to train ALBATROSS and STARLING notably differ from our training dataset Human–IDRome. In particular, their datasets contain synthetic as well as biological IDP sequences, and the conformational ensembles were generated via coarse-grained MD using an adapted version of the Mpipi force field [joseph2021] rather than CALVADOS-2. We therefore additionally retrained these models from scratch on the Human–IDRome dataset, and report results using both the pretrained and retrained versions of these models in Table 1.
@@ D.1 STARLING
Following the preprocessing in STARLING [novak2025], we first downsampled the frames for each sequence in our Human–IDRome dataset to reduce the correlation between conformers. We found that keeping every 20th frame was sufficient to stabilise model training, resulting in 50 conformers for each sequence. After downsampling, we used the same hyperparameters and methodology as in [novak2025] to sequentially retrain the STARLING VAE and DDPM models from scratch using our train and validation splits.
@@ D.2.2 Retraining
We use the same model architecture hyperparameters (number of hidden layers, hidden size) for each feature as used in the published V2 models. We found that we could improve training stability and performance by replacing the loss function used by [lotthammer2024] with the mean of the L1 loss over a batch rather than the sum, and performing a grid search over batch sizes $\{$64, 128, 256}\} and learning rates $\{$1e-3, 5e-3, 1e-2}\} for each model. We report the best test $R^{2}$ score achieved over the grid search for each feature in Table 1, and the hyperparameters used for each model in Table 3.
@@ D.2.3 $R^{2}$ score calculation
The $R^{2}$ scores attained in our evaluation of ALBATROSS are notably lower than those reported in the original ALBATROSS work [lotthammer2024]. This discrepancy can be partly explained by a difference in the definition of $R^{2}$ used between our work and theirs. In [lotthammer2024], the authors define the $R^{2}$ score as the square of the Pearson correlation coefficient between the true and predicted values, whereas here we define $R^{2}$ as the coefficient of determination.
@@ D.2.4 Flory scaling parameters
We compute the Flory scaling parameters by fitting a power-law relationship between the Euclidean distance of residue pairs and their sequence separation. Following the methodology of the Human-IDRome paper [tesei2024idrome], we exclude residue pairs with a sequence separation of less than five, as these short-range interactions are governed by local chain stiffness and deviate from the global scaling law, which may differ from that used by ALBATROSS [lotthammer2024].